OpenStudy (caerus):

linear equation of order one:

OpenStudy (caerus):

@jiteshmeghwal9

OpenStudy (caerus):

\[\int\limits_{}^{}-\frac{ v }{ e^v }dv\]

OpenStudy (jiteshmeghwal9):

wht d u want me to find ?

OpenStudy (jiteshmeghwal9):

@Caerus

OpenStudy (caerus):

how can i integrate this? ^.^

OpenStudy (jiteshmeghwal9):

\[\int\limits-ve^{-v} dv\]use by parts methd from here

OpenStudy (jiteshmeghwal9):

\[\text{Integration}=v \int\limits e^{-v}dv-\int\limits {\left( \frac{d(v)}{dv}*\int\limits e^{-v}dv \right)dv}\]

OpenStudy (caerus):

i got ve^-v + e^-v +c

OpenStudy (jiteshmeghwal9):

It should be \[ve^{-v}-e^{-v}+c\]

OpenStudy (caerus):

ok thanks

OpenStudy (caerus):

moving on to the next problem

OpenStudy (jiteshmeghwal9):

\[\text{integration}=-\left( v \int\limits\limits e^{-v}dv-\int\limits\limits {\left( \frac{d(v)}{dv}*\int\limits\limits e^{-v}dv \right)dv} \right)\]